1.Find the second and third derivatives in the simplest form of

y=x5-3x2-2x+5

2. Find by implicit differentiation of

z2-3zy+y2=6z-2y

1
Expert's answer
2021-09-20T16:11:02-0400

Q1. Given function "y=x^5-3x^2-2x+5\\"


"y'=\frac{dy}{dx}=5x^4-6x-2\\
y''=\frac{d^2y}{dx^2}=20x^3-6x-2\\
y'''=\frac{d^3y}{dx^3}=60x^2-6"


Q2. Given function "z^2-3zy+y^2=6z-2y"

"2z\frac{dy}{dz}-3y\frac{dy}{dz}-3z+2y=6\frac{dy}{dz}-2\\
\Rightarrow (2z-3y-6)\frac{dy}{dz}=3z-2y-2\\
\Rightarrow \frac{dy}{dz}=\frac{(3z-2y-2)}{(2z-3y-6)}"